### v2

# open-sci-ref-0.01: open and reproducible reference baselines for language model and dataset comparison

Abstract
We introduce open-sci-ref, a family of dense transformer models trained as research baselines across multiple model (0.13B to 1.7B parameters) and token scales (up to 1T) on 8 recent open reference datasets. Evaluating the models on various standardized benchmarks, our training runs set establishes reference points that enable researchers to assess the sanity and quality of alternative training approaches across scales and datasets. Intermediate checkpoints allow comparison and studying of the training dynamics. The established reference baselines allow training procedures to be compared through their scaling trends, aligning them on a common compute axis. Comparison of open reference datasets reveals that training on NemoTron-CC HQ consistently outperforms other reference datasets, followed by DCLM-baseline and FineWeb-Edu. In addition to intermediate training checkpoints, the release includes logs, code, and downstream evaluations to simplify reproduction, standardize comparison, and facilitate future research  [^1] Model weights and intermediate checkpoints are available at open-sci-ref-0.01 collection; code for reproducing training, evaluation and raw experiments data open-sci-ref-0.01 repository..
1 Introduction
To make guided progress in researching and improving foundation models and datasets, it is necessary to compare training procedures across scales. Yet reproducible and strong reference baselines for such comparisons are scarce, with existing ones relying on outdated recipes, older datasets, or being restricted to small scales. In this work, we provide a set of training runs and resulting base language models as improved reference baselines. These can be used to compare and examine other training procedures, assessing their sanity and quality across open reference datasets and scales. We rely on classical dense transformers implementations, as in works like Pythia biderman2023 , DCLM li2025 , OLMo(groeneveld-etal-2024-olmo, ), PolyPythias vanderwal2025 , DataDecide magnusson2025datadecide that use fully open, reproducible training pipelines and also provide intermediate checkpoints for studying training dynamics. Our reference baselines allow comparisons across four models scales (up to 1.7B parameters), trained on eight different datasets and up to larger 1T tokens scale, using a modern training procedure which ensures strong baselines for comparison.
2 Methods & Experiment Setup
2.1 Datasets, tokenization, and training details
We consider 8 datasets: 7 established references, namely C4 raffel2020exploring , Pile gao2020pile , SlimPajama cerebras2023slimpajama , FineWeb-Edu-1.4T (v1.0.0) penedo2024 , DCLM-baseline li2025 , Nemotron-CC-HQ su2025nemotroncc , HPLT-2.0 (english subset) burchell2025 , and CommonCorpus langlais2025 , a multi-lingual Common Crawl–based data pool not filtered for training. We tokenize datasets using the GPT-NeoX-20B tokenizer black2022gptneox , resulting in a vocabulary size of 50,304. For all datasets, we train reference dense transformer models with 130M, 400M, 1.3B, and 1.7B parameters using Megatron-LM megatron-lm . Following other strong reference models like Llama, Qwen, DCLM, and DeepSeek, we employ QK normalization naseer2021 in the model architecture for more stable training and constant learning rate with linear cooldown schedule (also known as WSD or trapezoid schedule (hu2024minicpm, ; hagele2024scaling, )), which facilitates training resumption and better supports scaling law derivation across token budgets.
2.2 Model architecture
We follow a custom design starting from the Llama (touvron2023llamaopenefficientfoundation, ) architecture. Different from standard Llama, we include biases in all linear layers (QKV projections, attention, and FFN), as removing them in ablation experiments at the 1.3B, 50BT scale resulted in performance deterioration, and we employ QK normalization (henry2020qknorm, ) for improved training stability and better performance. We use SwiGLU (shazeer2020glu, ) and tied embedding weights across all model scales. Tab. 1 summarizes the model configurations. Hereafter, we refer to this model design as open-sci-ref.
Params (B) (Non-Emb + Emb) Layers Hidden Heads FFN Hidden GPU Memory usage TFLOPS (6N)
0.1 + 0.03 = 0.13 22 512 8 2256 0.89GB $7.8\cdot 10^{8}$
0.35 + 0.05 = 0.40 22 1024 16 3840 2.88GB $2.4\cdot 10^{9}$
1.21 + 0.10 = 1.31 24 2048 32 5440 7.544GB $7.9\cdot 10^{9}$
1.61 + 0.10 = 1.71 24 2048 32 8192 9.884GB $1.0\cdot 10^{10}$
Table 1: open-sci-ref model architecture and scales.
2.3 Sanity checks and downstream evaluation
To validate the training procedure of our reference models, we perform a comparison with established baselines on the open datasets C4 and FineWeb-Edu, using open-sourced HuggingFace (HF) models with 1.7B parameters, trained on 350B tokens (denoted as HF-ref). We evaluate both HF-ref and open-sci-ref on standardized benchmarks (Tab. 2). Since open-sci-ref matches or outperforms HF-ref—despite the latter being trained on 50B more tokens—we conclude that our training procedure is well established and tuned for a given 1.7B model scale.
Model Dataset Tokens Params (B) Compute (FLOPS) Avg arc-c [10] arc-e [10] hellaswag [10] mmlu [5] copa [0] HF-ref-1.7B C4 350B 1.7 $3.57\cdot 10^{21}$ 0.506 0.32 0.66 0.64 0.25 0.74 open-sci-ref-1.7B C4 300B 1.7 $3.06\cdot 10^{21}$ 0.548 0.34 0.67 0.68 0.26 0.81 HF-ref-1.7B FineWeb-Edu 350B 1.7 $3.57\cdot 10^{21}$ 0.541 0.46 0.77 0.62 0.25 0.78 open-sci-ref-1.7B FineWeb-Edu 300B 1.7 $3.06\cdot 10^{21}$ 0.549 0.44 0.75 0.63 0.26 0.76
Table 2: Comparing HuggingFace (HF) and open-sci-ref reference baselines on C4 and FineWeb-Edu-1.4T using standardized evaluations. [n] provides few-shot numbers. open-sci-ref (300B tokens) matches or outperforms the results of HF (350B tokens) reference baseline.
Evaluation. Having establishing the soundness of our training procedure, we execute training of reference baselines and evaluate every intermediate checkpoint using LM-eval-harness eval-harness . We evaluate on Copa copa , Openbookqa mihaylov2018OpenBookQA , Lambada-openai paperno2016lambada , Winogrande sakaguchi2019winogrande on 0-shot setting, MMLU hendrycks2021measuring in 5-shot, and Commonsense-qa talmor2019commonsenseqa , Piqa bisk2020piqa , Arc-challenge clark2018thinkarx , Arc-easy, Hellaswag zellers2019hellaswag , Boolq clark2019boolq with 10-shot, resulting in 11 evaluations in total.
Control experiments. We perform additional experiments, testing the effect of different context lengths (4096, 8192, 16384 tokens) on the 1.7B model trained for 1T tokens, and compare various learning rate schedules (Suppl.  A.3). For additional hyperparameter tuning details, see Suppl. Sec. B
3 Results
3.1 Dataset comparison across scales
Figure 1: Comparison of average performance across 11 evaluation benchmarks for open-sci-ref models trained on 8 datasets with 300B tokens at different model scales. Dataset rankings remain consistent across scales, with differences becoming more pronounced at larger scales.
Figure 2: Downstream performance across tasks of open-sci-ref models with 1.7B parameters trained on 300B tokens over different datasets. While some tasks yield clear dataset rankings (e.g., ARC, Hellaswag, Lambada), others provide insufficient signal for meaningful dataset comparison.
Fig.˜1 reports the 11-task average performance across four model sizes, each trained on a different dataset at the 300B-token scale. Our results clearly show that recent state-of-the-art datasets such as Nemotron-CC-HQ and DCLM provide better performance, with gaps widening for larger models, while the ranking of top datasets remains stable across scales. By comparison, CommonCorpus, as an example of a multilingual data pool not filtered for base model training, unsurprisingly performs worse than English-dominant reference datasets filtered and curated for training.
Fig.˜2 shows task-level performance for the 1.7B model trained on 300B tokens on the reference datasets. We see pronounced differences between datasets on some evaluation tasks, though the overall ranking remains consistent. On MMLU, dataset comparison is difficult due to the lack of signal for all datasets except Nemotron-CC-HQ. This is likely because this dataset contains generated instruction-like samples with various rephrasings, boosting the ability to deal with multiple choice questions in MMLU, whereas other datasets yield scores close to random guessing (0.25). We additionally report results for the three top-performing datasets at the 1T-tokens scale (see Suppl. Sec. A, Tab. 3). For details on distributed training, see Suppl. Sec. C
3.2 Scaling trends based comparison
Fig.˜3 shows scaling behavior of our open-sci-ref reference baseline models when varying the training compute budget (increasing both model and token scale) for different datasets. By performing measurements across selected reference scales, a consistent dataset ranking can be established by looking at scaling trends across a broad span while aligning evaluated models on a common compute axis. The scaling trends can hint at promising strong datasets even given a low training budget setting and without deriving full scaling laws. Interestingly, our reference baseline 1.7B model trained on Nemotron-cc HQ for 1T token matches SmolLM2-1.7B (both 0.66 average score), achieving competitive results despite being trained with a simple single stage procedure on far fewer tokens (1T vs 11T), thus using roughly $11\times$ less compute (see also Tab. 5).
In addition to english dominant pre-trained baseline models, we take for comparison multi-lingually pretrained EuroLLM martins2024eurollm ; martins2025eurollm as it represents so far the strongest multi-lingual model coming from EU based efforts (significantly outpeforming similar motivated EU based models like Salamandra, Teuken or Occiglot). Well suited for comparison to our reference baselines, EuroLLM-1.7B used FineWeb-Edu as core part of its pre-training dataset, having at least 50% english data while adding multiple languages to reach total of 4T tokens, making it an interesting case for comparison.
It is important not to take the slopes of the scaling trends as indication how scalable the learning procedure is, without comparing the points to other references. For instance, for its model scale and invested compute, EuroLLM 1.7B strongly underperforms other learning procedures when properly taking into account pre-training compute. Eg, it has same performance level as SmolLM2 360M which has smaller model scale and used substantially less compute as evident from Fig.˜3. It underperforms 1.7B open-sci-ref models despite those using more than 10x less compute (see also Tab. 3). DCLM-1B model that has slightly less compute than EuroLLM 1.7B strongly outperforms it. Such points on smaller scales that are highly suboptimal can create the illusion of a learning procedure with strong scaling trend, as in the case of slope given for EuroLLM-1.7B - EuroLLM-9B line. Comparing EuroLLM-1.7B to other established references on common compute axis helps to avoid the illusion, revealing that steep slope towards EuroLLM-9B is due to the suboptimal performance on smaller 1.7B scale and not due to procedure’s scalability.
Figure 3: Scaling trends: open-sci-ref reference baselines (0.13, 0.4, 1.3, 1.7B model scales and 300B, 1T token scales) using 8 different datasets (solid lines) compared to other models (SmolLM2, Qwen 2.5, EuroLLM, DCLM) across various model and token scales (dashed lines), aligned on common compute axis (all tested models are base models). Evaluating reference baselines that use the same training procedure across scales while accounting for training compute provides dataset comparison and ranking. Nemotron-$C$-HQ, DCLM-baseline and FineWeb-Edu are strong reference datasets leading the ranking.
Tab. 3 shows more details on the per-task and average performance numbers for our reference baselines on 1.7B model and 300B, 1T token scales, providing a comparison to other models of similar sizes and with similar or different token budgets. We observe that EuroLLM-1.7B underperforms across all english benchmarks open-sci-ref-1.7B trained on 1T tokens of FineWeb-Edu data, and even on 300B tokens of FineWeb-Edu, open-sci-ref-1.7B still prevails on many benchmarks despite more than 10x less pre-training compute (more than 6x less compute on English, when comparing 300B vs. 2T tokens). This suggests that mixing languages can substantially weaken English benchmark performance and reduce problem-solving ability, as we see deterioration on English benchmarks despite substantially greater compute spent on English content.
Model Dataset Tokens Params Compute Avg copa lambada open wino mmlu arc-c arc-e boolq common hellaswag piqa
(B) (FLOPS) [0] [0] bookqa[0] grnd[0] [5] [10] [10] [10] sense[10] [10] [10]
gemma-2-2b 2.0T 2.60 $3.12\cdot 10^{22}$ 0.68 0.88 0.70 0.37 0.69 0.53 0.52 0.82 0.80 0.65 0.74 0.80
Qwen2.5-1.5B 18.0T 1.50 $1.62\cdot 10^{23}$ 0.67 0.83 0.62 0.36 0.63 0.61 0.52 0.81 0.78 0.76 0.68 0.77
DCLM-1B DCLM 4.0T 1.40 $3.36\cdot 10^{22}$ 0.66 0.90 0.67 0.43 0.68 0.47 0.48 0.78 0.75 0.62 0.74 0.79
SmolLM2-1.7B smolLM2 11.0T 1.70 $1.13\cdot 10^{23}$ 0.66 0.82 0.67 0.38 0.66 0.50 0.52 0.80 0.75 0.60 0.73 0.78
open-sci-ref-1.7B Nemotron 1T 1.70 $1.02\cdot 10^{22}$ 0.66 0.84 0.60 0.43 0.63 0.50 0.51 0.80 0.79 0.62 0.72 0.79
open-sci-ref-1.7B DCLM 1T 1.70 $1.02\cdot 10^{22}$ 0.57 0.79 0.68 0.40 0.64 0.24 0.44 0.76 0.69 0.19 0.70 0.77
open-sci-ref-1.7B FineWeb-Edu 1T 1.70 $1.02\cdot 10^{22}$ 0.56 0.81 0.54 0.43 0.63 0.26 0.47 0.76 0.67 0.20 0.67 0.76
open-sci-ref-1.7B FineWeb-Edu 300B 1.70 $3.06\cdot 10^{21}$ 0.55 0.76 0.52 0.42 0.61 0.26 0.44 0.75 0.67 0.19 0.63 0.76
HF-ref-1.7B FineWeb-Edu 350B 1.70 $3.57\cdot 10^{21}$ 0.54 0.78 0.50 0.37 0.58 0.25 0.46 0.77 0.66 0.19 0.62 0.75
EuroLLM-1.7B 4.0T 1.70 $4.08\cdot 10^{22}$ 0.52 0.74 0.53 0.33 0.59 0.27 0.39 0.73 0.61 0.19 0.60 0.74
Table 3: Performance of our 1.7B reference baselines trained on 1T tokens on Nemotron-$c$-hq, DCLM, FineWeb-Edu (also for 300B tokens) and several other baselines across various evals. Models are sorted by their average eval performance. HF-ref model is reference training run of 1.7B scale released by HuggingFace in the frame of their FineWeb study using 350B tokens. [n] for evaluation tasks indicate the number of shots. open-sci-ref 1.7B trained in Nemotron 1T tokens is a strong reference baseline, closely matching smolLM2 which uses 11x more compute, pointing to Nemotron-$c$-hq as strong reference training dataset.
4 Conclusion & Outlook
This work presents a research-based language model family open-sci-ref-0.01, providing a set of dense transformer models with up to 1.7B parameters trained on 50B, 300B, and 1T tokens on 8 estabilished reference datasets. These open baselines can be used to compare training procedures across scales, enabling easy sanity checks of training pipelines, and to validate novel models and datasets by mapping their evaluations against the reference baselines on a common compute axis.
Using the established reference baselines, we compare performance across dataset, obtaining dataset rankings that remain consistent across scales (Fig. 1,  3) and over the course of training (Fig. 1,  2). This consistency across scales and training durations provides stronger validation than any single-scale comparison, strengthening our analysis. We further compare against external models, aligning training procedures and reference baselines in a common frame of shared compute axis (Fig. 3). This reveals how weaker or stronger procedures can be identified using small scale baseline experiments. Nevertheless, establishing comparisons that retain validity across broad scales remains an open challenge and will require the derivation of full scaling laws nezhurina2025scaling ; we leave this to follow-up work.
Having derived reference baselines for comparison with any other learning procedures, in future work we will: (a) study permissively licensed datasets—CommonCorpus langlais2025commoncorpuslargestcollection , CommonPile kandpal2025commonpilev018tb , MixtureVitae mixturevitae2025dataset —and compare them to reference datasets which contain non-permissive content; (b) evaluate mixture-of-experts architectures against dense transformers; and (c) derive full scaling laws to enable robust and accurate comparisons at larger scales not covered here.
Acknowledgements
MN, JF, TC, NA, VK and JJ acknowledge co-funding by EU from Digital Europe Programme under grant no. 101195233 (openEuroLLM). MN and JJ acknowledge co-funding from EuroHPC Joint Undertaking programme under grant no. 101182737 (MINERVA), as well as funding by the Federal Ministry of Education and Research of Germany (BMBF) under grant no. 01IS24085C (OPENHAFM), under the grant 16HPC117K (MINERVA) and under the grant no. 01IS22094B (WestAI - AI Service Center West).
We gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this work by providing computing time through the John von Neumann Institute for Computing (NIC) on the supercomputer JUWELS Booster at Jülich Supercomputing Centre (JSC), EuroHPC Joint Undertaking for computing time and storage on the EuroHPC supercomputer LEONARDO, hosted by CINECA (Italy) and the LEONARDO consortium through an EuroHPC Extreme Access grant EHPC-EXT-2023E02-068, storage resources on JUST granted and operated by JSC and supported by Helmholtz Data Federation (HDF), computing time granted by the JARA and JSC on the supercomputer JURECA at JSC, and computing time granted on prototype JEDI via JUREAP (JUPITER Early Access Program) grant at JSC.
Further thanks go to support provided by supercomputing facilities and their teams, especially to Damian Alvarez and Mathis Bode from Juelich Supercomputer Center (JSC, Germany) and to Laura Morselli from CINECA (Italy).
We thank further for support & advising: Sampo Pyysalo, Guilherme Penedo, Quentin Anthony, Hynek Kydlíček, Mehdi Cherti, Tomer Porian, Adam Ibrahim, Jonathan Burdge, Aaron Klein, Ken Tsui, Harsh Raj, Huu Nguyen.
We also would like to express gratitude to all the people who are working on making code, models and data publicly available, advancing community based research and making research more reproducible. Specifically, we would like to thank Open-$\Psi$ (Open-Sci) Collective [^2] https://discord.gg/GsKh4mBVcv, openEuroLLM Consortium [^3] https://openeurollm.eu and all the members of the LAION Discord server [^4] https://discord.gg/BZqhreFazY community for providing fruitful ground for scientific exchange and open-source development.
Supplementary
Appendix A Additional experimental results
A.1 Results per tasks on 1T tokens
Fig.˜4 shows the performance while training the 1.7B model for the top 3 datasets while increasing the budget to 1T tokens. The ranking remains consistent, Nemotron staying mostly on the top, followed or matched by DCLM and FineWeb-Edu (with the exception of Lambada, where DCLM dominates), while performance is going up as expected when increasing token scale.
Figure 4: Downstream performance across tasks of open-sci-ref 1.7B models trained for 1T tokens on the top 3 reference datasets. Nemotron ranks highest across tasks, except on Lambada.
Aligning models on the common compute axis for evaluation as done in Fig. 3 and Tab. 3 can provide a good overview where various learning procedures stand relative to each other, which is often difficult if not making compute that went into the training explicit. Proper scaling trend estimation requires the models across scales to be close to compute optimal. In this work, measured reference baseline points are not guaranteed to be close to compute optimal. Measurements that are executed for scaling law derivation can identify models trained close to compute optimal Pareto front. Such measurements give much more accurate predictions about learning procedure behavior than scaling trends presented here. This however require more dense measurements and hyperparameter tuning effort [34].
A.2 Further insights from multilingual benchmarks
We perform comparisons on two parallel multilingual benchmarks: MMMLU, which contains translated MMLU instructions by OpenAI and Belebele [38].
Interestingly, for MMMLU, we see in Fig.˜5 that despite being pre-trained only on Nemotron-CC-HQ, which is a largely English-dominated dataset, open-sci-ref-0.01 shows stronger performance in language understanding problem solving on multi-lingual MMLU than EuroLLM-1.7B across multiple languages, which were explicitly included in the training set for EuroLLM. This demonstrates that heavily English-dominant pre-training enables language comprehension and problem solving across multiple languages if the dataset composition is strong as in the case with Nemotron-CC-HQ.
In Fig.˜6, we show the performance on Belebele while training a 1.7B model on Nemotron-CC-HQ, which echoes the surprising results seen on MMMLU. While the dataset is supposed to be English only, the model trained exhibits above random-chance performance (25%) on languages such as French, Spanish, Danish, or Portuguese. This is a surprising result which ought to be investigated. We hypothesize that it could come from the fact that the tolerance threshold used when filtering for English text lets further substantial data in another languages to slip into Nemontron dataset [10].
Figure 5: Performance of English pre-trained open-sci-ref on MMMLU, which is a version of the MMLU benchmark translated to multiple languages by OpenAI employing human professional translators.
Figure 6: Average performance on Belebele for different languages while training a 1.7B model on Nemotron-$C$-HQ.
A.3 Further reference baselines
Learning rate schedules. To make sure that reference training shows similar performance when executed with various schedules, we check training with a cosine schedule to match close constant learning rate and cooldown schedule (WSD). We execute baseline training on the C4 dataset with 50B tokens, using both the cosine and WSD schedules, tuning hyperparameters for each schedule. We see matching performance for cosine and WSD on downstream evaluations.
Context length. We also perform training with various context lengths while making sure that the global batch size remains constant across different variants. We train 2048, 4096, 8192, and 16384 context lengths, and observe no difference in downstream evaluations (Suppl. Fig.˜7). When varying context length, we also adapt the base for RoPE correspondingly (10k for 2048 and 4096, 100k as an alternative for 4096, 500k for 8192, and 1M for 16384).
Figure 7: Comparing training with different context length. Training run for open-sci-ref 1.7B model on 1T token of Nemotron-$C$-HQ, using various context length of 4096 (RoPE base 100k), 8192 (RoPE base 500k) and 16384 (RoPE base 1M). We observe no difference in average performance, ensuring models with larger context length have no drop in performance compared to reference context length of 4096.
Dropout. We also perform a test enabling or disabling dropout (hidden and attention), without observing any relevant difference in performance (the released models have dropout=0.1).
Appendix B Hyperparameter tuning
We perform hyperparameter tuning, measuring test loss on C4 and FineWeb-Edu, selecting hyperparameters that result in the lowest test loss on the datasets using 50B training runs with 1.3B model scale. We first experimented with a standard cosine learning rate schedule (Suppl. Tab. 5, tuning hyperparameters to match the HuggingFace reference models trained for 350B on C4 and FineWeb-Edu.
After achieving the match and thus checking the training procedure sanity, we switched to a constant learning + cooldown (WSD) schedule and tuned hyperparameters until we matched or outperformed results obtained with a cosine learning rate schedule on both 50B and 300B token budgets for a 1.3B model scale, which is still small enough for the tuning experiments to be affordable (Suppl. Tab. 4). The tuned WSD schedule training procedure was then used for all follow-up experiments and all model scales (0.13B, 0.4B, 1.3B, 1.7B). Tuned procedure uses weight decay of 0.05, global batch size of 4M tokens, cooldown of 20% of total training duration. Warmup and learning rate depend on the chosen total token budget. For larger token budgets >> 300B, a large warmup of 25 000 iterations turned out to give performance gains.
Tokens Global bs (tokens) Iters lr Warmup Cooldown (20%)
50B 2.65M 18839 $2\cdot 10^{-3}$ 6000 3767
50B 4.12M 11921 $4\cdot 10^{-3}$ 1000 2384
300B 2.09M 143052 $1\cdot 10^{-3}$ 5000 28610
300B 4.12M 72661 $4\cdot 10^{-3}$ 25000 14532
1T 4.12M 242204 $4\cdot 10^{-3}$ 25000 48440
Table 4: Hyperparameters for different token budgets, WSD lr schedule.
Tokens Global bs (tokens) iters lr warmup
50B 2.65M 18839 3e-3 6000
50B 4.03M 12406 6e-3 1000
300B 2.09M 143052 3e-3 5000
Table 5: Hyperparameters for cosine lr schedule
Appendix C Distributed training on supercomputers
We make use of various publicly funded European supercomputers to perform distributed training necessary for obtaining reference models. We conducted training experiments using Megatron-LM, employing NVIDIA containers for each machine. We used Leonardo (CINECA, 64GB A100), JUWELS Booster (JSC, 40GB A100), JEDI (JSC, early JUPITER prototype, 96GB H100), and JUPITER (JSC, 96GB H100). For each model scale, a combination of GPU numbers was selected that was providing good throughput and GPU utilization given the global batch size. We obtain good GPU utilization of 150-200 TFLOPS/s/GPU on A100 machines and 320-400 TFLOPS/s/GPU on H100 machines while working with hundreds of GPUs.
Model (B) GPUs micro bs context length global bs (smpl/token) TFLOPS/ GPU/s Tokens/ GPU/s Run Time, A100 h (50B/300B/1T) 0.13 84 12 4096 1008/4.12M 114 87710 1.89/11.31/37.70 0.4 100 10 4096 1000/4.09M 157 44550 3.12/18.71/62.35 1.3 252 4 4096 1008/4.12M 172 16800 3.28/19.68/65.60 1.7 216 6 2048 1296/2.65M 167 18490 3.48/20.87/69.56 1.7 252 4 4096 1008/4.12M 184 14490 3.80/22.81/76.05 1.7 252 2 8192 504/4.12M 186 12320 4.47/26.84/89.47 1.7 252 1 16384 252/4.12M 201 10080 5.47/32.80/109.32
Table 6: open-sci-ref reference baseline distributed training runs on Leonardo.
Model (B) GPUs micro bs context length global bs (smpl/token) TFLOPS/ GPU/s Tokens/ GPU/s Run Time,h (50B/300B/1T) GPU H100 h (50B/300B/1T)
1.3 64 16 2048 1024/2.09M 320 35370 6.13/36.81/122.70 393/2356/7853
1.7 128 2 16384 256/4.19M 395 19820 5.48/32.85/109.50 701/4205/14016
(a) executed on JEDI
Model (B) GPUs micro bs context length global bs (smpl/token) TFLOPS/ GPU/s Tokens/ GPU/s Run Time,h (50B/300B/1T) GPU H100 h (50B/300B/1T)
1.7 128 8 4096 1024/4.19M 370 29170 3.72/22.32/74.39 476/2856/9521
(b) executed on JUPITER
Model (B) GPUs micro bs context length global bs (smpl/token) TFLOPS/ GPU/s Tokens/ GPU/s Run Time,h (50B/300B/1T) GPU A100 h (50B/300B/1T)
1.3 328 6 2048 1968/4.03M 158 17400 2.43/14.60/48.66 798/4788/15961
(c) executed on JUWELS Booster
Table 7: open-sci-ref reference baseline distributed training runs
Appendix D Author contributions
  • Marianna Nezhurina: established major part of dataset processing (tokenization) and training infrastructure (Megatron-LM container based workflow), conducted scaling tests for distributed training, wrote routines for evaluation based on lm-eval-harness, downloaded and tokenized datasets, co-designed experiments, converted DCLM base models (1B, 7B) to HF and ran evaluation for the scaling plot. Co-wrote code for tables and figures.
  • Joerg Franke: downloaded and tokenized datasets, transferred datasets between various machines, co-designed experiments, conducted training experiments, co-wrote the paper
  • Taishi Nakamura: wrote checkpoint conversion routines from Megatron to HuggingFace format for custom open-sci-ref models to enable easy evaluation via lm-eval-harness.
  • Timur Carstensen: automated and performed conversion of all the Megatron checkpoints to HuggingFace format for evaluation using a script provided by Marianna and Taishi, helped running evaluations, provided the tooling to parse all hyperparameters from the logs, performed the evaluations and visualizations on MMMLU, co-wrote the paper
  • Niccolò Ajroldi: helped running evaluations and fixed a bug in lm-eval-harness to handle custom paths, composed figures, co-wrote the paper.
  • Ville Komulainen: uploaded all the intermediate and final checkpoints to HuggingFace, co-organized HuggingFace repository.
  • David Salinas: wrote the infrastructure to automate and perform large batch of evaluations, ran most of the evaluations, wrote code to generate most of the tables and figures, co-wrote the paper.
  • Jenia Jitsev: coordination and project supervision; acquired compute resources; designed the experiments (scales, architecture configuration, evaluation selection), wrote training scripts for various supercomputers, conducted major fraction of training runs, wrote routines for transferring datasets across supercomputers, downloaded and transferred the datasets across the machines, helped running evaluation, composed tables and figures, wrote the paper.