### v3

# open-sci-ref-0.01: open and reproducible reference baselines for language model and dataset comparison

Abstract
We introduce open-sci-ref, a family of dense transformer models trained as research baselines across multiple model (0.13B to 1.7B parameters) and token scales (up to 1T) on 8 recent open reference datasets. Evaluating the models on various standardized benchmarks, our training runs set establishes reference points that enable researchers to assess the sanity and quality of alternative training approaches across scales and datasets. Intermediate checkpoints allow comparison and studying of the training dynamics. The established reference baselines allow training procedures to be compared through their scaling trends, aligning them on a common compute axis. Comparison of open reference datasets reveals that training on NemoTron-CC HQ consistently outperforms other reference datasets, followed by DCLM-baseline and FineWeb-Edu. In addition to intermediate training checkpoints, the release includes logs, code, and downstream evaluations to simplify reproduction, standardize comparison, and facilitate future research  [^1] Model weights and intermediate checkpoints are available at https://huggingface.co/collections/open-sci/open-sci-ref-001; code for reproducing training, evaluation and raw experiments data https://github.com/LAION-AI/open-sci-ref-0.01.
1 Introduction
To make guided progress in researching and improving foundation models and datasets, it is necessary to compare training procedures across scales. Yet reproducible and strong reference baselines for such comparisons are scarce, with existing ones relying on outdated recipes, older datasets, or being restricted to small scales. In this work, we provide a set of training runs and resulting base language models as improved reference baselines. These can be used to compare and examine other training procedures, assessing their sanity and quality across open reference datasets and scales. We rely on classical dense transformers implementations, as in works like Pythia (Biderman et al. 2023), DCLM (Li et al. 2024), OLMo(Groeneveld et al. 2024), PolyPythias (van der Wal et al. 2025), DataDecide (Magnusson et al. 2025) that use fully open, reproducible training pipelines and also provide intermediate checkpoints for studying training dynamics. Our reference baselines allow comparisons across four models scales (up to 1.7B parameters), trained on eight different datasets and up to larger 1T tokens scale, using a modern training procedure which ensures strong baselines for comparison.
2 Methods & Experiment Setup
2.1 Datasets, tokenization, and training details
We consider 8 datasets: 7 established references, namely C4 (Raffel et al. 2020), Pile (Gao et al. 2020), SlimPajama (Soboleva et al. 2023), FineWeb-Edu-1.4T (v1.0.0) (Penedo et al. 2024), DCLM-baseline (Li et al. 2024), Nemotron-CC-HQ (Su et al. 2025), HPLT-2.0 (english subset) (Burchell et al. 2025), and CommonCorpus (Langlais et al. 2025), a multi-lingual Common Crawl–based data pool not filtered for training. We tokenize datasets using the GPT-NeoX-20B tokenizer (Black et al. 2022), resulting in a vocabulary size of 50,304. For all datasets, we train reference dense transformer models with 130M, 400M, 1.3B, and 1.7B parameters using Megatron-LM (Shoeybi et al. 2019). Following other strong reference models like Llama, Qwen, DCLM, and DeepSeek, we employ QK-Norm (Henry et al. 2020) in the model architecture for more stable training and constant learning rate with linear cooldown schedule (also known as WSD or trapezoid schedule (Hu et al. 2024; Hägele et al. 2024)), which allows resuming training from intermediate states easily and also saves compute cost when performing experiments across various token budgets, for instance as required for scaling law derivation.
2.2 Model architecture
We follow a custom design starting from the Llama (Touvron et al. 2023) architecture. Different from standard Llama, we include biases in all linear layers (QKV projections, attention, and FFN), as removing them in ablation experiments at the 1.3B, 50BT scale resulted in performance deterioration, and employ QK-Norm for improved training stability and better performance (Wortsman et al. 2023). We use SwiGLU (Shazeer 2020) and tied embedding weights across all model scales. Tab. 1 summarizes the model configurations. Hereafter, we refer to this model design as open-sci-ref.
Params (B) (Non-Emb + Emb) Layers Hidden Heads FFN Hidden GPU Memory usage TFLOPS (6N)
0.1 + 0.03 = 0.13 22 512 8 2256 0.89GB $7.8\cdot 10^{8}$
0.35 + 0.05 = 0.40 22 1024 16 3840 2.88GB $2.4\cdot 10^{9}$
1.21 + 0.10 = 1.31 24 2048 32 5440 7.544GB $7.9\cdot 10^{9}$
1.61 + 0.10 = 1.71 24 2048 32 8192 9.884GB $1.0\cdot 10^{10}$
Table 1: open-sci-ref model architecture and scales.
2.3 Hyperparameter tuning
Tokens Global bs (tokens) iters lr warmup
50B 2.65M 18839 3e-3 6000
50B 4.03M 12406 6e-3 1000
300B 2.09M 143052 3e-3 5000
Table 2: Hyperparameters for cosine lr schedule
Tokens Global bs (tokens) Iters lr Warmup Cooldown (20%)
50B 2.65M 18839 $2\cdot 10^{-3}$ 6000 3767
50B 4.12M 11921 $4\cdot 10^{-3}$ 1000 2384
300B 2.09M 143052 $1\cdot 10^{-3}$ 5000 28610
300B 4.12M 72661 $4\cdot 10^{-3}$ 25000 14532
1T 4.12M 242204 $4\cdot 10^{-3}$ 25000 48440
Table 3: Hyperparameters for different token budgets, WSD lr schedule.
We perform hyperparameter tuning by first measuring test loss on C4 and FineWeb-Edu, selecting hyperparameters that result in the lowest test loss on the datasets using 50B training runs with 1.3B model scale.
Learning rate schedules. We first tune hyperparams using standard cosine learning rate schedule (Tab. 2). We execute baseline training on the C4 dataset with 50B and 300B tokens, and obtain hyperparams for cosine learning schedule that correspond to minimal observed test loss. We then switch to constant learning + cooldown (WSD) and tune learning rate until we match or outperform minimal test loss observed with cosine schedule.
After achieving the match and thus checking WSD training procedure sanity, we switch to WSD schedule. The tuned WSD schedule training procedure (Tab. 3) is then used for all follow-up experiments and all model scales (0.13B, 0.4B, 1.3B, 1.7B). Tuned procedure uses weight decay of 0.05, global batch size of 4M tokens, cooldown of 20% of total training duration. Warmup and learning rate depend on the chosen total token budget. For larger token budgets >> 300B, a large warmup of 25 000 iterations turned out to give performance gains.
2.4 Sanity checks and downstream evaluation
To validate the tuned training procedure of our reference models, we perform a comparison with established baselines on the open datasets C4 and FineWeb-Edu, using open-source HuggingFace (HF) models with 1.7B parameters, trained on 350B tokens (denoted as HF-ref). We evaluate both HF-ref and open-sci-ref on standardized benchmarks (Tab. 4). Since open-sci-ref matches or outperforms HF-ref—despite the latter being trained on 50B more tokens—we conclude that our training procedure is well established and tuned for a given 1.7B model scale.
Model Dataset Tokens Params (B) Compute (FLOPS) Avg arc-c [10] arc-e [10] hellaswag [10] mmlu [5] copa [0] HF-ref-1.7B C4 350B 1.7 $3.57\cdot 10^{21}$ 0.506 0.32 0.66 0.64 0.25 0.74 open-sci-ref-1.7B C4 300B 1.7 $3.06\cdot 10^{21}$ 0.548 0.34 0.67 0.68 0.26 0.81 HF-ref-1.7B FineWeb-Edu 350B 1.7 $3.57\cdot 10^{21}$ 0.541 0.46 0.77 0.62 0.25 0.78 open-sci-ref-1.7B FineWeb-Edu 300B 1.7 $3.06\cdot 10^{21}$ 0.549 0.44 0.75 0.63 0.26 0.76
Table 4: Comparing HuggingFace (HF) and open-sci-ref reference baselines on C4 and FineWeb-Edu-1.4T using standardized evaluations. [n] provides few-shot numbers. open-sci-ref (300B tokens) matches or outperforms the results of HF (350B tokens) reference baseline.
Evaluation. Having establishing the soundness of our training procedure, we execute training of reference baselines and evaluate every intermediate checkpoint using LM-eval-harness (Gao et al. 2021). We evaluate on Copa (Roemmele, Bejan, and Gordon 2011), Openbookqa (Mihaylov et al. 2018), Lambada-openai (Paperno et al. 2016), Winogrande (Sakaguchi et al. 2021) on 0-shot setting, MMLU (Hendrycks et al. 2021) in 5-shot, and Commonsense-qa (Talmor et al. 2019), Piqa (Bisk et al. 2020), Arc-challenge (Clark et al. 2018), Arc-easy, Hellaswag (Zellers et al. 2019), Boolq (Clark et al. 2019) with 10-shot, resulting in 11 evaluations in total.
Control experiments. We perform additional experiments, testing the effect of different context lengths (4096, 8192, 16384 tokens) on the 1.7B model trained for 1T tokens (Suppl. Sec. A.1).
3 Results
3.1 Dataset comparison across scales
Reference baselines on 300B tokens. Fig.˜1 reports the 11-task average performance across four model sizes, each trained on a different dataset at the 300B-token scale. Our results clearly show that recent state-of-the-art datasets such as Nemotron-CC-HQ and DCLM provide better performance, with gaps widening for larger models, while the ranking of top datasets remains stable across scales. By comparison, CommonCorpus, as an example of a multilingual data pool not filtered for base model training, unsurprisingly performs worse than English-dominant reference datasets filtered and curated for training.
Fig.˜2 shows task-level performance for the 1.7B model trained on 300B tokens on the reference datasets. We see pronounced differences between datasets on some evaluation tasks, though the overall ranking remains consistent. On MMLU, dataset comparison is difficult due to the lack of signal for all datasets except Nemotron-CC-HQ. This is likely because this dataset contains generated instruction-like samples with various rephrasings, boosting the ability to deal with multiple choice questions in MMLU, whereas other datasets yield scores close to random guessing (0.25). For details on distributed training, see Suppl. Sec. B
Figure 1: Comparison of average performance across 11 evaluation benchmarks for open-sci-ref models trained on 8 datasets with 300B tokens at different model scales. Dataset rankings remain consistent across scales, with differences becoming more pronounced at larger scales.
Figure 2: Downstream performance across tasks of open-sci-ref models with 1.7B parameters trained on 300B tokens over different datasets. While some tasks yield clear dataset rankings (e.g., ARC, Hellaswag, Lambada), others provide insufficient signal for meaningful dataset comparison.
Reference baselines on 1T tokens. Fig.˜3 shows the performance of training the 1.7B model for the top 3 datasets while increasing the budget to 1T tokens. The ranking remains consistent, Nemotron staying mostly on the top, followed or matched by DCLM and FineWeb-Edu (with the exception of Lambada, where DCLM dominates), while performance is going up as expected when increasing token scale.
Figure 3: Downstream performance across tasks of open-sci-ref 1.7B models trained for 1T tokens on the top 3 reference datasets. Nemotron ranks highest across tasks, except on Lambada.
Tab. 5 shows the per-task and average performance numbers for our reference baselines on 1.7B model and 300B, 1T token scales, providing a comparison to other models of similar sizes and with similar or different token budgets. In addition to strong english dominant pre-trained baseline models like DCLM-1B, we take for comparison multi-lingually pretrained EuroLLM (Martins et al. 2024, 2025) as it represents so far the strongest multi-lingual model coming from EU based efforts (significantly outpeforming similar motivated EU based models like Salamandra, Teuken or Occiglot). Well suited for comparison to our reference baselines, EuroLLM-1.7B used FineWeb-Edu as core part of its pre-training dataset, having at least 50% english data while adding multiple languages to reach total of 4T tokens, making it an interesting case for comparison. We observe that EuroLLM-1.7B underperforms across all english benchmarks open-sci-ref-1.7B trained on 1T tokens of FineWeb-Edu data, and even on 300B tokens of FineWeb-Edu, open-sci-ref-1.7B still prevails on many benchmarks. This again shows that mixing languages can lead to strongly diminished problem solving performance, even if using substantially more compute for English (2x in case of EuroLLM).
Model Dataset Tokens Params Compute Avg copa lambada open wino mmlu arc-c arc-e boolq common hellaswag piqa
(B) (FLOPS) [0] [0] bookqa[0] grnd[0] [5] [10] [10] [10] sense[10] [10] [10]
gemma-2-2b 2.0T 2.60 $3.12\cdot 10^{22}$ 0.68 0.88 0.70 0.37 0.69 0.53 0.52 0.82 0.80 0.65 0.74 0.80
Qwen2.5-1.5B 18.0T 1.50 $1.62\cdot 10^{23}$ 0.67 0.83 0.62 0.36 0.63 0.61 0.52 0.81 0.78 0.76 0.68 0.77
DCLM-1B DCLM 4.0T 1.40 $3.36\cdot 10^{22}$ 0.66 0.90 0.67 0.43 0.68 0.47 0.48 0.78 0.75 0.62 0.74 0.79
SmolLM2-1.7B smolLM2 11.0T 1.70 $1.13\cdot 10^{23}$ 0.66 0.82 0.67 0.38 0.66 0.50 0.52 0.80 0.75 0.60 0.73 0.78
open-sci-ref-1.7B Nemotron 1T 1.70 $1.02\cdot 10^{22}$ 0.66 0.84 0.60 0.43 0.63 0.50 0.51 0.80 0.79 0.62 0.72 0.79
open-sci-ref-1.7B DCLM 1T 1.70 $1.02\cdot 10^{22}$ 0.57 0.79 0.68 0.40 0.64 0.24 0.44 0.76 0.69 0.19 0.70 0.77
open-sci-ref-1.7B FineWeb-Edu 1T 1.70 $1.02\cdot 10^{22}$ 0.56 0.81 0.54 0.43 0.63 0.26 0.47 0.76 0.67 0.20 0.67 0.76
open-sci-ref-1.7B FineWeb-Edu 300B 1.70 $3.06\cdot 10^{21}$ 0.55 0.76 0.52 0.42 0.61 0.26 0.44 0.75 0.67 0.19 0.63 0.76
HF-ref-1.7B FineWeb-Edu 350B 1.70 $3.57\cdot 10^{21}$ 0.54 0.78 0.50 0.37 0.58 0.25 0.46 0.77 0.66 0.19 0.62 0.75
EuroLLM-1.7B 4.0T 1.70 $4.08\cdot 10^{22}$ 0.52 0.74 0.53 0.33 0.59 0.27 0.39 0.73 0.61 0.19 0.60 0.74
Table 5: Performance of our 1.7B reference baselines trained on 1T tokens on Nemotron-$c$-hq, DCLM, FineWeb-Edu (also for 300B tokens) and several other baselines across various evals. Models are sorted by their average eval performance. HF-ref model is reference training run of 1.7B scale released by HuggingFace in the frame of their FineWeb study using 350B tokens. [n] for evaluation tasks indicate the number of shots. open-sci-ref 1.7B trained in Nemotron 1T tokens is a strong reference baseline, closely matching smolLM2 which uses 11x more compute, pointing to Nemotron-$c$-hq as strong reference training dataset.
Aligning models on the common compute axis for evaluation as done in Fig. 4 and Tab. 5 can provide a good overview where various learning procedures stand relative to each other, which is often difficult if not making compute that went into the training explicit. Again, for EuroLLM 1.7B we see that for its model scale and invested compute, it strongly underperforms other learning procedures. Eg, it has same performance level as SmolLM2 360M which has smaller model scale and used substantially less compute, also matching or underperforming 1.7B open-sci-ref models that used less compute. DCLM-1B model that has slightly less compute than EuroLLM 1.7B strongly outperforms it. It is therefore important to not take the slopes of the scaling trends as indication how scalable the learning procedure is, without comparing the points to other reference procedures. Points on smaller scales that are highly suboptimal can create the illusion of a strongly scalable learning procedure by making the slope very steep. Eg, slope as defined by EuroLLM-1.7B and EuroLLM-9B (Martins et al. 2025) performance is steep due to EuroLLM-1.7B being heavily suboptimal, as becomes apparent by comparison to reference baselines. Proper scaling trend estimation requires the models across scales to be close to compute optimal. Such measurements are executed for scaling law derivation, which identifies models trained close to compute optimal Pareto front and gives thus much more accurate predictions about learning procedure behavior than scaling trends presented here, however requiring more dense measurements and hyperparameter tuning effort.
3.2 Scaling trends based comparison
Figure 4: Scaling trends of open-sci-ref-0.01 baselines: 0.13B (300BT), 0.4B (300BT), 1.3B (300BT), 1.7B (300BT), and 1.7B (1TT) trained across 8 datasets (solid lines). Dashed lines show comparison with external models across parameter and token scales.
Fig.˜4 shows scaling behavior of our open-sci-ref reference baseline models when varying the training compute budget (increasing both model and token scale) for different datasets. By performing measurements across selected reference scales, a consistent dataset ranking can be established by looking at scaling trends across a broad span while aligning evaluated models on a common compute axis. The scaling trends can hint at promising strong datasets even given a low training budget setting and without deriving full scaling laws. Tab. 5 reports the per-task and average performance numbers, providing a comparison of our 1.7B reference baselines with other models of similar sizes and with similar or different token budgets. Interestingly, our reference baseline 1.7B model trained on Nemotron-cc HQ for 1T token matches SmolLM2-1.7B (both 0.66 average score), achieving competitive results despite being trained with a simple single stage procedure on far fewer tokens (1T vs 11T), thus using roughly $11\times$ less compute.
Drawing further on our reference baselines, we compare open-sci-ref-1.7B trained on FineWeb-Edu (300B and 1T tokens) with EuroLLM-1.7B (Martins et al. 2024), aiming to investigate the challenges of multilingual pretraining. EuroLLM-1.7B relied on FineWeb-Edu as the core of its pre-training corpus, containing at least 50% English and extended with multiple languages to reach a total of 4T tokens. We observe that EuroLLM-1.7B underperforms open-sci-ref-1.7B trained on 1T tokens of FineWeb-Edu across all English benchmarks; notably, even the 300B-token open-sci-ref-1.7B surpasses EuroLLM-1.7B on many tasks. This suggests that mixing languages can substantially weaken English benchmark performance and reduce problem-solving ability, despite EuroLLM’s much larger overall token budget (4T) and substantially greater compute spent on English content ($2\times$) compared with open-sci-ref-1.7B at 1T.
3.3 Further insights from multilingual benchmarks
Figure 5: Performance of English pre-trained open-sci-ref on MMMLU, which is a version of the MMLU benchmark translated to multiple languages by OpenAI employing human professional translators.
We perform comparisons on two parallel multilingual benchmarks: MMMLU, which contains translated MMLU instructions by OpenAI and Belebele (Bandarkar et al. 2024).
Interestingly, for MMMLU, we see in Fig.˜5 that despite being pre-trained only on Nemotron-CC-HQ, which is a largely English-dominated dataset, open-sci-ref-0.01 shows stronger performance in language understanding problem solving on multi-lingual MMLU than EuroLLM-1.7B across multiple languages, which were explicitly included in the training set for EuroLLM. This demonstrates that heavily English-dominant pre-training enables language comprehension and problem solving across multiple languages if the dataset composition is strong as in the case with Nemotron-CC-HQ.
Figure 6: Average performance on Belebele for different languages while training a 1.7B model on Nemotron-$C$-HQ.
In Fig.˜6, we show the performance on Belebele while training a 1.7B model on Nemotron-CC-HQ, which echoes the surprising results seen on MMMLU. While the dataset is supposed to be English only, the model trained exhibits above random-chance performance (25%) on languages such as French, Spanish, Danish, or Portuguese. This is a surprising result which ought to be investigated. We hypothesize that it could come from the fact that the tolerance threshold used when filtering for English text lets further substantial data in another languages to slip into Nemontron dataset (Su et al. 2025).
4 Conclusion & Outlook
This work presents a research-based language model family open-sci-ref-0.01, providing a set of dense transformer models with up to 1.7B parameters trained on 50B, 300B, and 1T tokens on 8 estabilished reference datasets. These open baselines can be used to compare training procedures across scales, enabling easy sanity checks of training pipelines, and to validate novel models and datasets by mapping their evaluations against the reference baselines on a common compute axis.
Using the established reference baselines, we compare performance across dataset, obtaining dataset rankings that remain consistent across scales (Fig. 1, 4) and over the course of training (Fig. 1, 2). This consistency across scales and training durations provides stronger validation than any single-scale comparison, strengthening our analysis. We further compare against external models, aligning training procedures and reference baselines in a common frame of shared compute axis (Fig. 4). This reveals how weaker or stronger procedures can be identified using small scale baseline experiments. Nevertheless, establishing comparisons that retain validity across broad scales remains an open challenge and will require the derivation of full scaling laws (Nezhurina et al. 2025); we leave this to follow-up work.
Having derived reference baselines for comparison with any other learning procedures, in future work we will: (a) study permissively licensed datasets—CommonCorpus (Langlais et al. 2025), CommonPile (Kandpal et al. 2025), MixtureVitae (Nguyen et al. 2025)—and compare them to reference datasets which contain non-permissive content; (b) evaluate mixture-of-experts architectures against dense transformers; and (c) derive full scaling laws to enable robust and accurate comparisons at larger scales not covered here.
Acknowledgements
MN, JF, TC, NA, VK and JJ acknowledge co-funding by EU from Digital Europe Programme under grant no. 101195233 (openEuroLLM). MN and JJ acknowledge co-funding from EuroHPC Joint Undertaking programme under grant no. 101182737 (MINERVA), as well as funding by the Federal Ministry of Education and Research of Germany (BMBF) under grant no. 01IS24085C (OPENHAFM), under the grant 16HPC117K (MINERVA) and under the grant no. 01IS22094B (WestAI - AI Service Center West).
We gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this work by providing computing time through the John von Neumann Institute for Computing (NIC) on the supercomputer JUWELS Booster at Jülich Supercomputing Centre (JSC), EuroHPC Joint Undertaking for computing time and storage on the EuroHPC supercomputer LEONARDO, hosted by CINECA (Italy) and the LEONARDO consortium through an EuroHPC Extreme Access grant EHPC-EXT-2023E02-068, storage resources on JUST granted and operated by JSC and supported by Helmholtz Data Federation (HDF), computing time granted by the JARA and JSC on the supercomputer JURECA at JSC, and computing time granted on prototype JEDI via JUREAP (JUPITER Early Access Program) grant at JSC.
Further thanks go to support provided by supercomputing facilities and their teams, especially to Damian Alvarez and Mathis Bode from Juelich Supercomputer Center (JSC, Germany) and to Laura Morselli from CINECA (Italy).
We thank further for support & advising: Sampo Pyysalo, Guilherme Penedo, Quentin Anthony, Hynek Kydlíček, Mehdi Cherti, Tomer Porian, Adam Ibrahim, Jonathan Burdge, Aaron Klein, Ken Tsui, Harsh Raj, Huu Nguyen.
We also would like to express gratitude to all the people who are working on making code, models and data publicly available, advancing community based research and making research more reproducible. Specifically, we would like to thank Open-$\Psi$ (Open-Sci) Collective [^2] https://discord.gg/GsKh4mBVcv, openEuroLLM Consortium [^3] https://openeurollm.eu and all the members of the LAION Discord server [^4] https://discord.gg/BZqhreFazY community for providing fruitful ground for scientific exchange and open-source development.
Supplementary
Appendix A Additional experimental results
A.1 Further reference baselines
Context length. We also perform training with various context lengths while making sure that the global batch size remains constant across different variants. We train 2048, 4096, 8192, and 16384 context lengths, and observe no difference in downstream evaluations (Suppl. Fig.˜7). When varying context length, we also adapt the base for RoPE correspondingly (10k for 2048 and 4096, 100k as an alternative for 4096, 500k for 8192, and 1M for 16384).
Figure 7: Comparing training with different context length. Training run for open-sci-ref 1.7B model on 1T token of Nemotron-$C$-HQ, using various context length of 4096 (RoPE base 100k), 8192 (RoPE base 500k) and 16384 (RoPE base 1M). We observe no difference in average performance, ensuring models with larger context length have no drop in performance compared to reference context length of 4096.
Dropout. We also perform a test enabling or disabling dropout (hidden and attention), without observing any relevant difference in performance (the released models have dropout=0.1).
Appendix B Distributed training on supercomputers
We make use of various publicly funded European supercomputers to perform distributed training necessary for obtaining reference models. We conducted training experiments using Megatron-LM, employing NVIDIA containers for each machine. We used Leonardo (CINECA, 64GB A100), JUWELS Booster (JSC, 40GB A100), JEDI (JSC, early JUPITER prototype, 96GB H100), and JUPITER (JSC, 96GB H100). For each model scale, a combination of GPU numbers was selected that was providing good throughput and GPU utilization given the global batch size. We obtain good GPU utilization of 150-200 TFLOPS/s/GPU on A100 machines and 320-400 TFLOPS/s/GPU on H100 machines while working with hundreds of GPUs.
Model (B) GPUs micro bs context length global bs (smpl/token) TFLOPS/ GPU/s Tokens/ GPU/s Run Time, h (50B/300B/1T) GPU A100 h (50B/300B/1T) 0.13 84 12 4096 1008/4.12M 114 87710 1.89/11.31/37.70 158/950/3167 0.4 100 10 4096 1000/4.09M 157 44550 3.12/18.71/62.35 312/1871/6235 1.3 252 4 4096 1008/4.12M 172 16800 3.28/19.68/65.60 827/4959/16530 1.7 216 6 2048 1296/2.65M 167 18490 3.48/20.87/69.56 751/4508/15026 1.7 252 4 4096 1008/4.12M 184 14490 3.80/22.81/76.05 958/5749/19164 1.7 252 2 8192 504/4.12M 186 12320 4.47/26.84/89.47 1127/6764/22546 1.7 252 1 16384 252/4.12M 201 10080 5.47/32.80/109.32 1377/8265/27549
Table 6: open-sci-ref reference baseline distributed training runs on Leonardo.
Model (B) GPUs micro bs context length global bs (smpl/token) TFLOPS/ GPU/s Tokens/ GPU/s Run Time,h (50B/300B/1T) GPU H100 h (50B/300B/1T)
1.3 64 16 2048 1024/2.09M 320 35370 6.13/36.81/122.70 393/2356/7853
1.7 128 2 16384 256/4.19M 395 19820 5.48/32.85/109.50 701/4205/14016
(a) executed on JEDI
Model (B) GPUs micro bs context length global bs (smpl/token) TFLOPS/ GPU/s Tokens/ GPU/s Run Time,h (50B/300B/1T) GPU H100 h (50B/300B/1T)
1.7 128 8 4096 1024/4.19M 370 29170 3.72/22.32/74.39 476/2856/9521
(b) executed on JUPITER
Model (B) GPUs micro bs context length global bs (smpl/token) TFLOPS/ GPU/s Tokens/ GPU/s Run Time,h (50B/300B/1T) GPU A100 h (50B/300B/1T)
1.3 328 6 2048 1968/4.03M 158 17400 2.43/14.60/48.66 798/4788/15961
(c) executed on JUWELS Booster
Table 7: open-sci-ref reference baseline distributed training runs
Appendix C Author contributions
  • Marianna Nezhurina: established major part of dataset processing (tokenization) and training infrastructure (Megatron-LM container based workflow), conducted scaling tests for distributed training, wrote routines for evaluation based on lm-eval-harness, downloaded and tokenized datasets, co-designed experiments, converted DCLM base models (1B, 7B) to HF and ran evaluation for the scaling plot. Co-wrote code for tables and figures.
  • Joerg Franke: downloaded and tokenized datasets, transferred datasets between various machines, co-designed experiments, conducted training experiments, co-wrote the paper
  • Taishi Nakamura: wrote checkpoint conversion routines from Megatron to HuggingFace format for custom open-sci-ref models to enable easy evaluation via lm-eval-harness.
  • Timur Carstensen: automated and performed conversion of all the Megatron checkpoints to HuggingFace format for evaluation using a script provided by Marianna and Taishi, helped running evaluations, provided the tooling to parse all hyperparameters from the logs, performed the evaluations and visualizations on MMMLU, co-wrote the paper
  • Niccolò Ajroldi: helped running evaluations and fixed a bug in lm-eval-harness to handle custom paths, composed figures, co-wrote the paper.
  • Ville Komulainen: uploaded all the intermediate and final checkpoints to HuggingFace, co-organized HuggingFace repository.
  • David Salinas: wrote the infrastructure to automate and perform large batch of evaluations, ran most of the evaluations, wrote code to generate most of the tables and figures, co-wrote the paper.
  • Jenia Jitsev: coordination and project supervision; acquired compute resources; designed the experiments (scales, architecture configuration, evaluation selection), wrote training scripts for various supercomputers, conducted major fraction of training runs, wrote routines for transferring datasets across supercomputers, downloaded and transferred the datasets across the machines, helped running evaluation, composed tables and figures, wrote the paper.